OpenStudy (anonymous):

Find the sum of the geometric series in which a1=972, n=5 and r=-1/3

OpenStudy (anonymous):

The formula for the sum is a(1-r^n)/(1-r) so 972(1-(-1/3)^5)/(1- 1/3) = 972(1+ 1/243) * 3 /2 = 972 * 244/243 * 3/2 = 12*122 = 1464